OpenStudy (anonymous):

the sum of the square of two consecutive integers is 41. Find the integers.

OpenStudy (anonymous):

Consecutive integers means numbers that are next to each other The consecutive integers of 14 and 15 make 29 Does this make sense? It is the basis of the problem.

OpenStudy (anonymous):

I think it does.

OpenStudy (anonymous):

x^2 + (x+1)^2 = 41

OpenStudy (anonymous):

Because of this, you can estimate what the answer to your question is Because it is 2 numbers, you can divide 41 by 2 41/2 = 20.5 No you have to pick 2 consecutive numbers that have an average of 20.5

OpenStudy (anonymous):

What integers are inbetween 20.5? 20 and 21 Does that make sense? 20+21 = 41

OpenStudy (anonymous):

it says the sum of SQUARE of two consecutive numbers

OpenStudy (anonymous):

Oh okay my bad. You just have to do it in reverse then. 20^2 = 400 21^2 = 441 441 - 400 = 41

OpenStudy (anonymous):

okay thank you KirbyLegs! That makes more sense

OpenStudy (anonymous):

Welcome :)

OpenStudy (anonymous):

no @KirbyLegs is wrong solve this x^2 + (x+1)^2 = 41

OpenStudy (anonymous):

then u'll get the ans

OpenStudy (anonymous):

it's not the difference. It's the sum

OpenStudy (anonymous):

yep